A grocery store used to sell bulk chocolate cookies for $0.65 for 100g. The price increased by $0.65 four 100g. What Is the new price for 400g of bulk chocolate chip cookies


1
Expert's answer
2021-09-27T16:36:50-0400

New price of 100g chocolate"=0.65+0.65=\$1.30"

Cost of 400g of bulk chocolate chips"=4\times 1.30=\$5.20"